OpenStudy (anonymous):

Show the steps to simplfiying sin(a-b)sin(a+b)=(cosb)^2-(cosa)^2

OpenStudy (anonymous):

This question is hard. D:

OpenStudy (anonymous):

If you get bored of this question I got 2 others.

OpenStudy (mertsj):

sin(a-b)sin(a+b)= [sin(a)cos(b)-cos(a)sin(b)][sin(a)cos(b)+cos(a)sin(b)]= sin^2(a)cos^2(b)-cos^2(a)sin^2(b)= (1-cos^2a)cos^2b-cos^2a(1-cos^2b)= cos^2b-cos^2acos^2b-cos^2a+cos^2acos^2b= cos^2b-cos^2a

OpenStudy (anonymous):

ohhhh, see i was going on the same path as you. But somehow I didn't follow through enough like you.